OpenStudy (anonymous):

Prove the following: || F1||= (||F3|| cos*beta)/(sin(alpha +beta))

OpenStudy (anonymous):

\[Prove the following: |\left| F _{1} \right|| = (|\left| F _{3} \right||\cos \beta) \div (\sin(\alpha +\beta))\]